Fermi National Laboratory
 

Accelerator Update for April 4, 2007 - April 6, 2007

 

Wednesday April 4

The day shift began with the Tevatron (TeV) recovering from a quench, with Operations monitoring the Recycler stash (83.5E10), with the Antiproton Source (Pbar) stacking (25.8E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

Operations ramped the TeV at 10:40 AM.

Operations put the TeV through a dry squeeze at 10:58 AM.

TeV experts began a study at 11:17 AM.

At 11:57 AM, Operations reported that a TeV quadrupole power supply (T:D0Q5) had tripped off during the first dry squeeze. Experts investigated.

Operations put the TeV through another dry squeeze at 12:01 PM so experts could investigate the DC lead flows. They determined that the lead flow was too cold and that the TeV should stay in a dry squeeze until the leads warmed up.

Operations established beam to the returning Meson MTest experiment (MT964) at 12:48 PM. T964 wants to understand the various chamber characteristics of a Gas Electron Monitor (GEM) developed at CERN. GEMs are x-ray sensitive beam monitors. The International Linear Collider (ILC) intends to use them as gap monitors.

A Pbar expert began a test of the Pbar target sweeping magnet at 1:20 PM.

A MI expert began a MI-8 collimator study at 2:29 PM.

TeV experts put the TeV through a wet squeeze at 2:33 PM.

Around 3:10 PM, the I- Source vacuum gauge tripped off. It reset.

A Pbar expert began a stacktail study at 3:42 PM.

The Pbar studies ended at 4:05 PM and Pbar resumed stacking.

Operations reported at 4:20 PM that the Recycler stashed antiprotons three times during the day shift.

Operations reported at 5:17 PM that they were turning on MiniBooNE power supplies. MiniBooNE hopes to resume taking beam on April 12, 2007.

MTest experimenters reported at 7:06 PM that they were done with the beam for the night.

The MI studies ended at 7:31 PM.

TeV experts completed a study. Operations loaded a 12x0 proton only store at 9 PM and experts checked the beam positions.

Operations put the TeV into shot setup at 9:35 PM.

Operations established store 5328 at 11:26 PM with an initial luminosity of 198.82E30.

Operations reported at midnight that the Recycler had stashed two times during the evening shift.

Thursday April 5

The midnight shift began with Operations monitoring store 5328 (167.46E30) and the stash (56.7E10), with Pbar stacking (19.4E10), and with NuMI taking beam.

At 1:22 AM, the I- Source vacuum gauge tripped off. It reset.

Operations established beam to MTest at 6:10 AM.

Operations reported at 7:53 AM that the Recycler had stashed three times during the midnight shift.

Per instructions from a MiniBooNE expert, operators turned off the MiniBooNE power supplies at 8:03 AM.

At 8:17 AM, a MI RF station (MIRF14) tripped off for the first of two times in six minutes. It reset.

At 8:22 AM, the I- Source vacuum gauge tripped off for the first of three times in about 90 minutes. It reset.

At 8:23 AM, MIRF14 tripped off. It reset.

At 9:07 AM, BRF6 tripped off for the first of ten times in two hours. After the last trip an expert investigated and reduced the power supply’s output voltage.

A MI expert began a multibatch slipstacking study at 9:36 AM.

Pbar experts began a stacktail study around 1:48 PM. The study lasted, off and on, about two and a half hours.

At 4:23 PM, the MI electrical safety system dropped out on a “door open” indication. This caused the loss of the Recycler stash. There was no open door. Interlocks technicians found and replaced a bad relay at MI-8.

Operations began the MI turn on at 5:42 PM.

Operations established beam to Pbar and NuMI at 5:52 PM.

The MTest experiment reported at 5:58 PM that they were through with the beam for the night.

At 6:30 PM, a Booster corrector magnet power supply (B:CEPSW1) tripped off and wouldn’t reset. Operations called in an expert. The expert replaced a power supply. Booster resumed sending beam at 8:25 PM.

At 10:16 PM, LRF2 tripped for the first of multiple times in about 23 minutes. It reset.

At 11:06 PM, a D0 cryo system cold compressor tripped off. Cryo system techs investigated.

At 11:07 PM, a Pbar Debuncher RF station (DRF1-5) tripped off. Operators investigated.

At 11:16 PM, LRF3 tripped off. It reset.

Operations reported at midnight that the Recycler had stashed once during the evening shift.

Friday April 6

The midnight shift began with Operations monitoring store 5328 (34.28), the stash (60.5E10), and the stack (42.5E10), with stacking off due to problems with DRF1-5, and with NuMI taking beam.

Pbar resumed stacking at 12:38 AM. Operators tried several things to fix DRF1-5 with no success. They bypassed the station.

Operations terminated store 5328 at 1:58 AM.

A TeV expert began a study period at 2:46 AM.

At 5:12 AM, the I- Source vacuum gauge had tripped off for the first of three times in two hours. It reset.

At 5:19 AM, Operations reported that MIRF2 had tripped off twice. Operators lowered the MI beam intensity.

Operations reported at 7:52 AM that the Recycler had stashed three times during the midnight shift.

At 7:56 AM, cryo system techs at CHL reported that their nitrogen plant was down. Experts investigated.

The Plans for Friday and the Weekend

The plans for today are for TeV experts to prepare the TeV for the next store and to conduct sextupole studies while the stash gets built up.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/06/2007    email Fermilab